Hubbard's Pet Foods is financed 80% by common stock and 20% by bonds. The expected return on the common stock is 14% and the rate of interest on the bonds is 6%. Assume that the bonds are default-free and that there are no taxes. Now assume that Hubbard's issues more debt and uses the proceeds to retire equity. The new financing mix is 25% equity and 75% debt.
If the debt is still default-free, what happens to the expected rate of return on equity? (Round your answer to 2 decimal places.)
  Expected rate of return on equity   %
What happens to the expected return on the package of common stock and bonds? (Round your answer to 2 decimal places.)
  Expected return on the package of common stock and bonds    %
Answer #1

Answer 1:

Current cost capital = rA = rE * E/V + rD * D/V = 14% * 80% + 6% * 20% = 12.40%

The new financing mix is 25% equity and 75% debt.

Expected rate of return on equity = rA + (D /E * (rA - rD) = 12.40% + (75% / 25% * (12.40% - 6%)) = 31.60%

Expected rate of return on equity = 31.60 %

Answer 2:

Expected return on the package of common stock and bonds = 31.60% * 25% + 6% * 75% = 12.40%

Expected return on the package of common stock and bonds = 12.40%